OpenStudy (anonymous):

Let m be a positive integer. Show that a mod m = b mod m if a ≡ b (mod m). Use only the definitions of mod m and congruence modulo m; do not use any theorems involving mod or congruence.

ganeshie8 (ganeshie8):

use the definition

ganeshie8 (ganeshie8):

\(a \equiv b \mod m\) means \(m | (a-b)\)

ganeshie8 (ganeshie8):

\(\large \implies a-b = mk\)

ganeshie8 (ganeshie8):

\(\large a = mk +b\) substitute this value in the left side of congruence you wanted to prove

ganeshie8 (ganeshie8):

\(\large a \mod m = (mk + b) \mod m \equiv b \mod m\) QED.

OpenStudy (anonymous):

Hare Ganesh!

ganeshie8 (ganeshie8):

let me know if something is not clear :)